OpenStudy (anonymous):

the sum of a number and -14 is -32 find the number

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

Your "sum of a number and -14" looks like -14 + x "is -32" looks like = -32 Together, this is -14 + x = -32 Then you use algebra to solve for x.

OpenStudy (anonymous):

so the equal sign is in the middle?

OpenStudy (anonymous):

Yes. You always have something = something else

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

youve helped me so much. ok so this should be pretty easy

OpenStudy (anonymous):

it is -18

OpenStudy (anonymous):

thats really easy

OpenStudy (anonymous):

six less than a number is -10 find the number

OpenStudy (anonymous):

Ideas?

OpenStudy (anonymous):

x-6=(-10)

OpenStudy (anonymous):

does thaat look right?

OpenStudy (anonymous):

Great :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that -6 im confused

OpenStudy (anonymous):

Yes. "6 less than a number" is the same as "a number subtract 6."

OpenStudy (anonymous):

+4?

OpenStudy (anonymous):

Let's test it. 4-6 = -2, so +4 won't work. x - 6 = -10 We want x by itself, so we need to get rid of that -6. To get rid of a debt of 6, we add 6. Then, what we do to one side, we do to the other. In summary, we add 6 to both sides. x = -4

OpenStudy (anonymous):

-4 I mean

OpenStudy (anonymous):

thats why its good to always check ure answers

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

twice the sum of a number and -2 is 16 find the number

OpenStudy (anonymous):

Ideas?

OpenStudy (anonymous):

2x+(-2)=-16

OpenStudy (anonymous):

Very close. You're taking twice the sum. So you do the sum part, and then you double it. 2[x + (-2)] = 16

OpenStudy (anonymous):

is that what the [ ] means?

OpenStudy (anonymous):

Oh, it's just another way to write parentheses. 2(x + (-2)) = 16

OpenStudy (anonymous):

so if I did not put parethsis I would have got it wrong?

OpenStudy (anonymous):

Yes, you need parentheses, because the -2 needs to be multiplied by 2 as well.